What is internal energy of a chemical reaction? Explain clearly using an equation.
Answer #1

Internal energy \left ( \Delta U \right ) of a chemical reaction is the heat of reaction \left ( Q_V \right ) at constant volume.

AU = Qv

Internal energy \left ( \Delta U \right ) of a chemical reaction is also calculated by the relation

AU= AH - RTA

Consider the following reaction

C(s) + O2(g) + CO2(9)

For the above reaction An=1-1=0

AU= AH - RTA

AU = -393.5 kJ/mol – 8.314 – X 298 K xol

AU = -393.5 kJ/mol